# SPA-Bench: A Comprehensive Benchmark for SmartPhone Agent Evaluation

### v1

@@ Abstract
Smartphone agents are increasingly important for helping users control devices efficiently, with (Multimodal) Large Language Model (MLLM)-based approaches emerging as key contenders. Fairly comparing these agents is essential but challenging, requiring a varied task scope, the integration of agents with different implementations, and a generalisable evaluation pipeline to assess their strengths and weaknesses. In this paper, we present SPA-Bench, a comprehensive SmartPhone Agent Benchmark designed to evaluate (M)LLM-based agents in an interactive environment that simulates real-world conditions. SPA-Bench offers three key contributions: (1) A diverse set of tasks covering system and third-party apps in both English and Chinese, focusing on features commonly used in daily routines; (2) A plug-and-play framework enabling real-time agent interaction with Android devices, integrating over ten agents with the flexibility to add more; (3) A novel evaluation pipeline that automatically assesses agent performance across multiple dimensions, encompassing seven metrics related to task completion and resource consumption. Our extensive experiments across tasks and agents reveal challenges like interpreting mobile user interfaces, action grounding, memory retention, and execution costs. We propose future research directions to ease these difficulties, moving closer to real-world smartphone agent applications.
@@ 1 Introduction
The growing capabilities of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) have broadened the application of AI agents across various domains Gur et al. (2023); Gou et al. (2023); Cai et al. (2023); Li et al. (2023a); Wang et al. (2023); Wu et al. (2023a). One promising area is smartphone control, where agents assist users in tasks like booking hotels or setting alarms. These agents, often powered by off-the-shelf Yang et al. (2023b); Wen et al. (2024); Wang et al. (2024b; a); Rawles et al. (2024a) or fine-tuned Zhan & Zhang (2023); Hong et al. (2024); Bai et al. (2024); Lu et al. (2024) (M)LLMs, rely heavily on these models as the “brains” for decision-making. The information these agents use to interact with smartphones can vary, with common methods involving direct screen observation Wang et al. (2024b; a); Zhan & Zhang (2023); Hong et al. (2024); Bai et al. (2024); Lu et al. (2024), accessing non-visible data via Android View Hierarchy or Extensible Markup Language (XML) Wen et al. (2024), or a combination of both Yang et al. (2023b); Rawles et al. (2024a).
@@
As the number of (M)LLM-based agents grows, fair performance comparison becomes crucial, leading to an increasing need for benchmarking them Chan et al. (2023); Liu et al. (2023a; b); Wu et al. (2023b). Regarding smartphone agent benchmarks, existing studies use three main approaches to evaluate agents: actions-based Xing et al. (2024), states-based Rawles et al. (2024a); Zhang et al. (2024); Lee et al. (2024); Wang et al. (2024c), or a hybrid of both Wang et al. (2024c). Each method faces specific difficulties: action-based evaluation may involve multiple correct sequences, while state-based methods struggle to determine the appropriate post-action state. A hybrid approach could mitigate these limitations, but the challenge lies in effectively utilising both action and state information.
@@
Despite these efforts, current research Rawles et al. (2024a); Xing et al. (2024); Zhang et al. (2024); Lee et al. (2024); Wang et al. (2024c) still has several key limitations: (1) The focus remains primarily on system and Google suite applications (apps) in English, which are often free from distractions like ads and pop-ups that could introduce complexity and randomness; (2) The number of evaluated agents is typically fewer than five, with some studies including only similar variants of the same agent; (3) Automated success detection methods frequently require human intervention (e.g., handcrafted validation logic for each task) or rely on data that may be inaccessible in certain cases (e.g., Android View Hierarchy data, which is unavailable in WebView apps Xing et al. (2024)).
@@
In this paper, we introduce SPA-Bench, a SmartPhone Agent Benchmark designed to evaluate more than 10 smartphone control agents in daily tasks. As illustrated in Figure 1, SPA-Bench comprises 340 tasks, including 150 single-app tasks and 20 cross-app tasks, in both English and Chinese apps, as well as third-party ones. It integrates 11 agents into a unified framework, connected to an automated evaluation pipeline that measures agent performance, with the ability to automatically expand to additional tasks beyond this benchmark. Our experiments show that agents using proprietary (M)LLMs outperform those using fine-tuned or open-source (M)LLMs, though they remain impractical for real-world deployment due to time and cost constraints. We also provide a detailed discussion on the challenges and future directions for smartphone agents, covering topics such as building perceptive mobile interfaces, reasoning mechanisms, and user-friendly applications.
@@
In summary, our comprehensive benchmark makes several key contributions: (1) a diverse task collection of 340 tasks with increasing difficulty, accompanied by human trajectory annotations. It covers both English and Chinese apps, including single-app and cross-app scenarios, and featuring 58 third-party apps (Section 3); (2) a plug-and-play agent framework supporting 11 agents, which allows for easy integration of new agents with minimal adaptation and offers features like automatic Android setup and multi-device emulator support (Section 4); (3) an automated evaluation pipeline assesses agent performance using task completion and resource consumption metrics. It employs success detection methods that achieve average F1 scores of 90.5% for single-app tasks and 84.5% for cross-app tasks compared to human evaluators (Section 5); and (4) extensive experiments across agents and tasks, providing a detailed analysis of current smartphone agent capabilities and limitations, while also offering directions for future research (Section 6).
@@ 2 Related Work
Smartphone Agent. Smartphone agents aim to automate tasks on mobile apps in a human-like way. Early agents, like Siri and Google Assistant, relied on system-level APIs and customisation, limiting their generality. Recently, (M)LLM-based agents have emerged, using the user interface (UI) to achieve a more general approach. These agents, with (M)LLMs as their “brains”, also require “hands” (actions) and “eyes” (observations) to interact with smartphones. They are based on either off-the-shelf or fine-tuned models and perform human-like actions (e.g., tapping, typing, and swiping). According to how they observe the UI, recent works are categorised into text-based, vision-based, and combined approaches. Text-based methods Wen et al. (2024); Rawles et al. (2024a) rely on UI document data (e.g., XML) or convert visual information into text, vision-based methods Wang et al. (2024b; a); Zhan & Zhang (2023); Hong et al. (2024); Bai et al. (2024); Lu et al. (2024) use screenshots to capture the complete visual context, while combined approaches Yang et al. (2023b); Rawles et al. (2024a) integrate both text and vision inputs for greater informativeness. SPA-Bench evaluates all three types of agents to provide a comprehensive comparison of their capabilities.
@@
Smartphone Agent Evaluation. Effective evaluation of smartphone agents is crucial for identifying limitations and guiding improvements. Success rate, which measures task completion, is the most commonly used metric, with some studies also considering efficiency. Success detection methods are generally classified into two types: human detection Yang et al. (2023b); Wang et al. (2024b; a), which is accurate but resource-intensive, and automated detection, which is less costly but varies in accuracy. Current automated methods primarily rely on hand-crafted validation logic, making them unscalable without human intervention. They are restricted to evaluating tasks involving apps that are limited to English-only and simpler apps (e.g., system, Google Suite, and open-source apps), with minimal coverage of other third-party ones. These automated methods can be further divided into action-based, state-based, and hybrid approaches. Action-based methods Xing et al. (2024) compare agents’ actions to human demonstrations but struggle with the non-unique nature of correct action sequences. State-based methods Rawles et al. (2024a); Zhang et al. (2024); Lee et al. (2024) assess whether essential states are reached but may miss minor actions. Hybrid approaches Wang et al. (2024c) combine state and action data for more accurate success detection. SPA-Bench introduces two hybrid approaches for evaluating single-app and cross-app tasks. Compared to other automated methods, our approaches support a wider range of apps and tasks. They do not rely on hand-crafted validation logic, making them adaptable without human intervention. Table 1 presents a comparison between our work and other automated evaluation-based smartphone agent benchmarks, highlighting our comprehensive evaluation of various agents in diverse tasks across multiple dimensions.
@@
Dataset Third-party app? Cross- app? Chinese app? Difficulty level? Number of tasks Number of agents Number of metrics Free of hand- crafted validation? Information for success detection
AndroidArena Xing et al. (2024) 221 1 4 Action only
AndroidWorld Rawles et al. (2024a) 116 3 1 State only
LlamaTouch Zhang et al. (2024) 495 4 1 State only
B-MoCA Lee et al. (2024) 60 3 1 State only
MobileAgentBench Wang et al. (2024c) 100 5 6 Action and State
@@ 3.1 Overview
The single-app tasks are grouped into three difficulty levels. In general, Level 1 requires fewer than five actions, Level 2 under ten actions, and Level 3 typically fewer than fifteen. Each set of tasks contains three tasks at varying difficulty levels, often sharing similar instructions but using different entities to prevent influence from earlier tasks (e.g., creating folders with different names). Examples of single-app tasks are shown in Figure 2.
Figure 2: A sample set of tasks within the Deliveroo app, annotated by human. In this example, simpler tasks form the foundation for more complex ones, resulting in shared trajectories in the initial stages. The final screenshots for tasks of all three difficulty levels are highlighted in corresponding colours. Each final screenshot highlights the key components used in coarse detection (explained further in Section 5.2), with the zoomed-in versions available in Appendix B.3.
For cross-app tasks, we refer to the recent work GUI Odyssey Lu et al. (2024), which defines six task types: General Tool, Information Management, Web Shopping, Media Entertainment, Social Sharing, and Multi-Apps. Our cross-app tasks include three Level 1 tasks for each of the first five types (requiring two apps) and five Level 2 tasks for the Multi-Apps type (requiring three apps). Appendix B.4 provides examples.
@@ 3.2 Task Construction
Our tasks were primarily constructed by human annotators. For single-app tasks, we selected commonly used apps and supplemented them with apps from related works Yang et al. (2023b); Wang et al. (2024b). Based on each app’s core features, tasks were created following an annotation guideline specifying: (1) A clear task description that reflects the task’s goal and difficulty level. For descriptions inspired by prior works, we standardised and assigned difficulty levels accordingly. (2) A human-executed trajectory, presented as a series of screenshots that avoid shortcuts and irrelevant actions. Between any two adjacent screenshots, only one action (e.g., tap, swipe, type) is allowed. The total number of actions in the human execution serves as the “golden steps” in our experiments. (3) Key components of the final state, which are pieces of text that must appear in the final screenshot if the task is successfully completed. We focus only on the final state because there may be multiple correct paths to complete the task, but they typically converge to the same final state Wang et al. (2024c). These key components are designed for future use, as detailed in Section 5.2.
For cross-app tasks, annotations include only task descriptions and human-executed trajectories due to the flexibility of final states. Most cross-app English tasks were drawn from GUI Odyssey Lu et al. (2024), and we reformatted descriptions and recollected trajectories where necessary.
@@ 4.1 A Unified Plug-and-play Framework
Our framework facilitates the execution of autonomous smartphone agents and tasks. As shown in Figure 3, the worker machine manages communication, providing task descriptions and receiving outcomes (trajectories and logs). It hosts multiple worker processes, each connecting an Android emulator [^1] https://developer.android.com/studio/run/emulator and an agent. Each agent interacts with the Android device by performing actions based on observations, such as taking screenshots and generating actions like taps, swipes, or long presses. The snapshot state is restored at the start of each experimental cycle.
@@
The framework is highly scalable. Unlike existing research Rawles et al. (2024a); Xing et al. (2024); Zhang et al. (2024); Lee et al. (2024); Wang et al. (2024c), which integrates a limited number of agents tightly into the framework, ours allows easy addition of new agents with minimal integration, ensuring each agent operates independently within an isolated environment. Details about the agents integrated into our framework are provided in Appendix C.
@@ 4.2 Snapshot-Based Emulator for Consistent Testing
The framework integrates Android emulators [^2] Certain English tasks involving WhatsApp and OneNote, and most Chinese tasks, were run only on physical Android devices instead of emulators due to strict app control measures. While the emulator can be replaced with physical Android devices, this would remove the snapshot feature mentioned here. as a scalable alternative to physical devices, replicating most Android functions for parallel testing and rapid experiment deployment. For instance, a 24-core CPU with 64GB RAM can support up to eight emulators or worker processes simultaneously, depending on the agents’ resource needs.
@@ 5.1 Metrics
Completion-related Metrics. (1) Success signal – a binary indicator of task success. For single-app and cross-app tasks, we develop two different hybrid approaches that leverage both action and state information, allowing for multiple valid execution paths. These approaches eliminate the need for human evaluators and handcrafted evaluation logic (details are provided in Section 5.2). (2) Step ratio – measures execution efficiency by comparing agent steps with human steps (the “golden steps” from Section 3.2). This is considered only when the task is successful (i.e., success signal is “true”). A higher ratio indicates more unnecessary actions and lower efficiency. (3) Termination reason – explains why the task was terminated, including reasons like self-reported completion, reaching the maximum step limit, or execution errors (e.g., invalid actions). (4) Premature termination signal – a binary indicator applicable only when the termination reason is self-reported completion. It is set to “true” when the success signal is “false”, indicating that the agent incorrectly believed the task had been successfully completed, such that it stopped too early. (5) Overdue termination signal – a binary indicator applicable only when the termination reason is reaching the maximum step limit. It is set to “true” when the success signal is “true”, meaning the agent mistakenly thought the task was uncompleted, resulting in unnecessary steps.

### v2

@@ Abstract
Smartphone agents are increasingly important for helping users control devices efficiently, with (Multimodal) Large Language Model (MLLM)-based approaches emerging as key contenders. Fairly comparing these agents is essential but challenging, requiring a varied task scope, the integration of agents with different implementations, and a generalisable evaluation pipeline to assess their strengths and weaknesses. In this paper, we present SPA-Bench, a comprehensive SmartPhone Agent Benchmark designed to evaluate (M)LLM-based agents in an interactive environment that simulates real-world conditions. SPA-Bench offers three key contributions: (1) A diverse set of tasks covering system and third-party apps in both English and Chinese, focusing on features commonly used in daily routines; (2) A plug-and-play framework enabling real-time agent interaction with Android devices, integrating over ten agents with the flexibility to add more; (3) A novel evaluation pipeline that automatically assesses agent performance across multiple dimensions, encompassing seven metrics related to task completion and resource consumption. Our extensive experiments across tasks and agents reveal challenges like interpreting mobile user interfaces, action grounding, memory retention, and execution costs. We propose future research directions to ease these difficulties, moving closer to real-world smartphone agent applications. SPA-Bench is available at https://ai-agents-2030.github.io/SPA-Bench/.
footnotetext: * Equal contribution.footnotetext: Corresponding authors: shaorui@hit.edu.cn, shaokun2@huawei.com.
@@ 1 Introduction
The growing capabilities of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs) have broadened the application of AI agents across various domains (Gur et al., 2023; Gou et al., 2023; Cai et al., 2023; Li et al., 2023a; Wang et al., 2023; Wu et al., 2023a). One promising area is smartphone control, where agents assist users in tasks like booking hotels or setting alarms. These agents can be broadly categorised into two main types: (1) agent-as-a-model (Lai et al., 2024), where fine-tuned or pre-trained (M)LLMs are customised for agentic tasks (Zhan & Zhang, 2023; Hong et al., 2024; Bai et al., 2024; Lu et al., 2024; Christianos et al., 2024; Wang et al., 2024d), and (2) agentic workflow (Shang et al., 2024), which typically relies on off-the-shelf models and modular designs to support agentic functionality (Yang et al., 2023b; Wen et al., 2024; Wang et al., 2024b; a; Rawles et al., 2024a). In both cases, these models act as the “brains” for decision-making. The information these agents use to interact with smartphones can vary, with common methods involving direct screen observation (Wang et al., 2024b; a; Zhan & Zhang, 2023; Hong et al., 2024; Bai et al., 2024; Lu et al., 2024), accessing non-visible data via Android View Hierarchy or Extensible Markup Language (XML) (Wen et al., 2024), or a combination of both (Yang et al., 2023b; Rawles et al., 2024a).
@@
As the number of (M)LLM-based smartphone agents grows, fair performance comparisons become crucial for identifying their strengths and shortcomings, leading to an increasing need for benchmarking (Chan et al., 2023; Liu et al., 2023a; b; Wu et al., 2023b). Regarding smartphone agent benchmarks, existing studies use three main approaches to evaluate agents: actions-based (Xing et al., 2024), states-based (Rawles et al., 2024a; Zhang et al., 2024; Lee et al., 2024; Wang et al., 2024c), or a hybrid of both (Wang et al., 2024c). Each method faces specific difficulties: action-based evaluation may involve multiple correct sequences, while state-based methods struggle to determine the appropriate post-action state. A hybrid approach could mitigate these limitations, but the challenge lies in effectively utilising both action and state information.
@@
Despite these efforts, current research (Rawles et al., 2024a; Xing et al., 2024; Zhang et al., 2024; Lee et al., 2024; Wang et al., 2024c) still has several key limitations: (1) The focus remains primarily on system and Google suite applications (apps) in English, which are often free from distractions like ads and pop-ups that could introduce complexity and randomness; (2) The number of evaluated agents is typically fewer than five, with some studies including only similar variants of the same agent; (3) Automated success detection methods frequently require human intervention (e.g., handcrafted validation logic for each task) or rely on data that may be inaccessible in certain cases (e.g., Android View Hierarchy data, which is unavailable in WebView apps (Xing et al., 2024)).
@@
In this paper, we introduce SPA-Bench, a SmartPhone Agent Benchmark designed to evaluate more than 10 smartphone control agents in daily tasks. As illustrated in Figure 1, SPA-Bench comprises 340 tasks, including 150 single-app tasks and 20 cross-app tasks, in both English and Chinese apps, as well as third-party ones. It integrates 11 agents into a unified framework based on their original implementations. This framework is linked to an automated evaluation pipeline that measures agent performance and can be applied to additional tasks beyond this benchmark, without requiring human inputs. Our experiments show that agents following the agentic workflow outperform those in the agent-as-a-model category, although the former one remain impractical for real-world deployment due to time and cost constraints. We also provide a detailed discussion on the challenges and future directions for smartphone agents, covering topics such as building perceptive mobile interfaces, reasoning mechanisms, and user-friendly applications.
@@
In summary, our comprehensive benchmark makes several key contributions: (1) a diverse task collection of 340 tasks with increasing difficulty, accompanied by human trajectory annotations. It covers both English and Chinese apps, including single-app and cross-app scenarios, and featuring 58 third-party apps (Section 3); (2) a plug-and-play agent framework supporting 11 agents, which allows for easy integration of new agents with minimal adaptation and offers features like automatic Android setup and multi-device emulator support (Section 4); (3) an automated and scalable evaluation pipeline assesses agent performance using task completion and resource consumption metrics. It employs success detection methods that achieve average F1 scores of 90.5% for single-app tasks and 84.5% for cross-app tasks compared to human evaluators (Section 5); and (4) extensive experiments across agents and tasks, providing a detailed analysis of current smartphone agent capabilities and limitations, while also offering directions for future research (Section 6).
@@ 2 Related Work
Smartphone Agent. Smartphone agents aim to automate tasks on mobile apps in a human-like way. Early agents, like Siri and Google Assistant, relied on system-level APIs and customisation, limiting their generality. Recently, (M)LLM-based agents have emerged, using the user interface (UI) to achieve a more general approach. These agents, with (M)LLMs as their “brains”, also require “hands” (actions) and “eyes” (observations) to interact with smartphones. They are based on either off-the-shelf or fine-tuned models and perform human-like actions (e.g., tapping, typing, and swiping). According to how they observe the UI, recent works are categorised into text-based, vision-based, and combined approaches. Text-based methods (Wen et al., 2024; Rawles et al., 2024a) rely on UI document data (e.g., XML) or convert visual information into text, vision-based methods (Wang et al., 2024b; a; Zhan & Zhang, 2023; Hong et al., 2024; Bai et al., 2024; Lu et al., 2024) use screenshots to capture the complete visual context, while combined approaches (Yang et al., 2023b; Rawles et al., 2024a) integrate both text and vision inputs for greater informativeness. SPA-Bench evaluates all three types of agents to provide a comprehensive comparison of their capabilities.
@@
Smartphone Agent Evaluation. Effective evaluation of smartphone agents is crucial for identifying limitations and guiding improvements. Success rate, which measures task completion, is the most commonly used metric, with some studies also considering efficiency. Success detection methods are generally classified into two types: human detection (Yang et al., 2023b; Wang et al., 2024b; a), which is accurate but resource-intensive, and automated detection, which is less costly but varies in accuracy. Current automated methods primarily rely on hand-crafted validation logic, making them unscalable without human intervention. They are restricted to evaluating tasks involving apps that are limited to English-only and simpler apps (e.g., system, Google Suite, and open-source apps), with minimal coverage of other third-party ones. These automated methods can be further divided into action-based, state-based, and hybrid approaches. Action-based methods (Xing et al., 2024) compare agents’ actions to human demonstrations but struggle with the non-unique nature of correct action sequences. State-based methods (Rawles et al., 2024a; Zhang et al., 2024; Lee et al., 2024) assess whether essential states are reached but may miss minor actions. Hybrid approaches (Wang et al., 2024c) combine state and action data for more accurate success detection. SPA-Bench introduces two hybrid approaches for evaluating single-app and cross-app tasks. Compared to other automated methods, our approaches support a wider range of apps and tasks. They do not rely on hand-crafted validation logic, making them adaptable without human intervention. Table 1 presents a comparison between our work and other automated evaluation-based smartphone agent benchmarks, highlighting our comprehensive evaluation of various agents in diverse tasks across multiple dimensions.
@@
Benchmark Third-party app? Cross- app? Chinese app? Difficulty level? Number of tasks Number of agents Number of metrics Free of hand- crafted validation? Information for success detection
AndroidArena  (Xing et al., 2024) 221 1 4 Action only
AndroidWorld  (Rawles et al., 2024a) 116 3 1 State only
LlamaTouch  (Zhang et al., 2024) 495 4 1 State only
B-MoCA  (Lee et al., 2024) 60 3 1 State only
MobileAgentBench  (Wang et al., 2024c) 100 5 6 Action and State
@@ 3.1 Overview
Figure 2: A sample set of tasks within the Deliveroo app, annotated by human. In this example, simpler tasks form the foundation for more complex ones, resulting in shared trajectories in the initial stages. The final screenshots for tasks of all three difficulty levels are highlighted in corresponding colours. Each final screenshot highlights the key components used in coarse detection (explained further in Section 5), with the zoomed-in versions available in Appendix B.3.
The single-app tasks are grouped into sets, with progressively increasing complexity across three difficulty levels. In each set, Level 1 tasks serve as foundational and straightforward activities, while Level 2 and Level 3 tasks introduce more complex requirements, such as handling intricate UI elements or animations. Generally, Level 1 tasks require fewer than five actions, while Level 2 tasks typically involve fewer than ten, and Level 3 tasks fewer than fifteen. While each set follows similar instructions, the tasks are designed to use distinct entities (e.g., creating folders with different names) to prevent any influence from earlier tasks. Examples of single-app tasks are shown in Figure 2. For cross-app tasks, we refer to the recent work GUI Odyssey (Lu et al., 2024), which defines six task types: General Tool, Information Management, Web Shopping, Media Entertainment, Social Sharing, and Multi-Apps. Unlike single-app tasks, cross-app difficulty levels are determined by the number of apps involved in a task. Level 1 tasks require interactions between two apps, while Level 2 tasks necessitate switching between three. As the number of apps increases, complexity arises not only from additional steps but also from inter-app dependencies and coordination. Our cross-app tasks include three Level 1 tasks for each of the first five types, and five Level 2 tasks for the Multi-Apps type. Appendix B.4 provides examples.
@@ 3.2 Task Construction
Our tasks were primarily constructed by human annotators. For single-app tasks, we selected commonly used apps and supplemented them with apps from related works (Yang et al., 2023b; Wang et al., 2024b). Based on each app’s core features, tasks were created following an annotation guideline specifying: (1) A clear task description that reflects the task’s goal and difficulty level. For descriptions inspired by prior works, we standardised them and assigned appropriate difficulty levels. (2) A human-executed trajectory presented as a series of screenshots. Between any two adjacent screenshots, only one action (e.g., tap, type) is allowed. The total number of actions in the human execution serves as the “golden steps” in our experiments. To ensure a reproducible and unbiased baseline, we instruct human annotators to avoid irrelevant actions and refrain from using shortcuts that are inherently dynamic, influenced by factors such as recommendation algorithms or user-specific history. (3) Key components of the final state, which are pieces of text that must appear in the final screenshot if the task is successfully completed. We focus only on the final state because there may be multiple correct paths to complete the task, but they typically converge to the same final state (Wang et al., 2024c). These key components are designed for future use, as detailed in Section 5.2. For cross-app tasks, annotations include only task descriptions and human-executed trajectories due to the flexibility of final states. Most cross-app English tasks were drawn from GUI Odyssey (Lu et al., 2024), and we reformatted descriptions and recollected trajectories where necessary.
@@ 4.1 A Unified Plug-and-play Framework
Our framework facilitates the execution of autonomous smartphone agents and tasks. As shown in Figure 3, the worker machine manages communication, providing task descriptions and receiving outcomes (trajectories and logs). It hosts multiple worker processes, each connecting an Android emulator and an agent. Each agent interacts with the Android device by performing actions based on observations, such as taking screenshots and generating actions like taps or swipes. The snapshot state is restored at the start of each experimental cycle. The framework is highly scalable. Unlike existing research (Rawles et al., 2024a; Xing et al., 2024; Zhang et al., 2024; Lee et al., 2024; Wang et al., 2024c), which integrates a limited number of agents tightly into the framework, ours allows easy addition of new agents with minimal integration, ensuring each agent operates independently within an isolated environment. Details about the agents integrated into our framework are provided in Appendix C.
@@ 4.2 Snapshot-Based Emulator for Consistent Testing
The framework integrates Android emulators as a scalable alternative to physical devices, replicating most Android functions for parallel testing and rapid experiment deployment. For instance, a 24-core CPU with 64GB RAM can support up to eight emulators or worker processes simultaneously, depending on the agents’ resource needs.
@@ 5.1 Metrics
Completion-related Metrics. (1) Success signal – a binary indicator of task success. For single-app and cross-app tasks, we develop two different hybrid approaches that leverage both action and state information, allowing for multiple valid execution paths. These approaches eliminate the need for human evaluators and handcrafted evaluation logic (details are provided in Section 5.2). (2) Step ratio – measures execution efficiency by comparing agent steps with human steps (the “golden steps” from Section 3.2). This is considered only when the task is successful (i.e., success signal is “true”). A higher ratio indicates more unnecessary actions and lower efficiency. (3) Termination reason – explains why the task was terminated, including self-reported completion (i.e., an agent proactively terminates a task based on its belief that the task has been completed successfully), reaching the configured maximum step limit, or execution errors (e.g., invalid actions).(4) Premature termination signal – a binary indicator applicable only when the termination reason is self-reported completion. It is set to “true” when the success signal is “false”, indicating the agent mistakenly believed the task was completed. This premature stopping reduces success rates by causing the agent to assume success before finishing the task. (5) Overdue termination signal – a binary indicator applicable only when the termination reason is reaching the maximum step limit. It is set to “true” when the success signal is “true”, meaning the agent mistakenly thought the task was incomplete. This results in unnecessary steps, reducing efficiency as the agent takes extra actions before concluding the task.